The dimensions of a rectangle were originally 10 units by 12 units. The area of the rectangle increased by 135 square units, and the dimensions were increased by the same amount. Find the dimensions of the new rectangle.
Answer by Boreal(15235)

(10+x)(12+x)=255
120+22x+x^2=255
x^2+22x-135=0
(x+27)(x-5)=0
x=5, only positive root
the new dimensions are 15 x 17 units or 255 units ^2, 135 units more than 120.